WATER REACTIVITY WITH TUNGSTEN OXIDES: FORMATION OF H_{2} FROM W_{2}O_{y}^{−} + H_{2}O REAC-TIONS

\underline{NICHOLAS J. MAYHALL}, DAVID W. ROTHGEB, EKRAM HOSSAIN, CAROLINE CHICK JAR-ROLD, AND KRISHNAN RAGHAVACHARI, Department of Chemistry, Indiana University, 800 E. Kirkwood Avenue, Bloomington, Indiana 47405, USA.

In a recent mass spectrometry/photoelectron spectroscopy study on the reactions between W_{2}O_{y}^{−} (y = 2–6) and water, Jarrold and coworkers (J. Chem. Phys., 130, 12431 2009) observed interesting differences in the reactivity of the different cluster ions. Particularly noteworthy is the observation that the only product with the incorporation of hydrogens is a single peak corresponding to W_{2}O_{6}H_{2}^{−}. As reactions between metal oxide clusters and small molecules such as water have high potential for catalytic applications, we aim to obtain a mechanistic understanding of this observed reactivity. Using electronic structure calculations, we have identified and characterized multiple modes of reactivity between unsaturated tungsten oxide clusters (W_{2}O_{y}^{−} (y = 4–6)) and water. By calculating the free energy corrected reaction profiles, our results provide an explanation for the formation of W_{2}O_{6}H_{2}^{−}. We propose a mechanism in which water reacts with a metal oxide cluster and eliminates H_{2}. The results from our calculations show that this is nearly a barrierless process for all sub-oxide clusters with the exception of W_{2}O_{5}^{−}.